# PaCo-RL: Advancing Reinforcement Learning for Consistent Image Generation with Pairwise Reward Modeling

### v1

@@ Abstract
Consistent image generation requires faithfully preserving identities, styles, and logical coherence across multiple images, which is essential for applications such as storytelling and character design. Supervised training approaches struggle with this task due to the lack of large-scale datasets capturing visual consistency and the complexity of modeling human perceptual preferences. In this paper, we argue that reinforcement learning (RL) offers a promising alternative by enabling models to learn complex and subjective visual criteria in a data-free manner. To achieve this, we introduce PaCo-RL, a comprehensive framework that combines a specialized consistency reward model with an efficient RL algorithm. The first component, PaCo-Reward, is a pairwise consistency evaluator trained on a large-scale dataset constructed via automated sub-figure pairing. It evaluates consistency through a generative, autoregressive scoring mechanism enhanced by task-aware instructions and CoT reasons. The second component, PaCo-GRPO, leverages a novel resolution-decoupled optimization strategy to substantially reduce RL cost, alongside a log-tamed multi-reward aggregation mechanism that ensures balanced and stable reward optimization. Extensive experiments across the two representative subtasks show that PaCo-Reward significantly improves alignment with human perceptions of visual consistency, and PaCo-GRPO achieves state-of-the-art consistency performance with improved training efficiency and stability. Together, these results highlight the promise of PaCo-RL as a practical and scalable solution for consistent image generation. The project page is available at https://x-gengroup.github.io/HomePage_PaCo-RL/.
@@ 1 Introduction
Image generation has made remarkable progress in producing high-quality and diverse images from textual prompts [baldridge2024imagen, saharia2022photorealistic, betker2023improving, esser2021taming, 4o_image_generation, wu2025qwenimagetechnicalreport, jia2025chatgen, xia2025idealrealunifieddataefficient, jia2024ssmgspatialsemanticmapguided]. However, achieving consistent image generation remains both challenging and crucial for various applications such as storytelling, advertising, and character creation [mao2024storyadaptertrainingfreeiterativeframework, tewel2024trainingfreeconsistenttexttoimagegeneration, chen2025interleavedscenegraphsinterleaved, zhou2024storydiffusion, liu2025onepromptonestory, huang2024incontextloradiffusiontransformers]. In this work, we study consistent image generation through two representative tasks: Image Editing [luo2025editscoreunlockingonlinerl] and Text-to-ImageSet [jia2025settleonetexttoimagesetgeneration], where the objective is to either align generated images with given references (Fig.˜1, left) or produce a coherent set of images from a single prompt (Fig.˜1, right). This is a challenging problem for supervised training approaches, due to the lack of large-scale datasets for visual consistency and the complexity of modeling human perceptions of consistency [wu2023humanpreferencescorev2, kirstain2023pickapicopendatasetuser, zhang2024learning, chen2025adiee, lin2024evaluating].
Recently, reinforcement learning (RL) has demonstrated strong potential in advancing image generation models [liu2025flowgrpotrainingflowmatching, xue2025dancegrpounleashinggrpovisual, zheng2025diffusionnftonlinediffusionreinforcement], offering a promising pathway to address the above challenges. Instead of relying on explicit supervision from carefully curated datasets, RL algorithms optimize generative models using feedback from reward models (RMs), which are often trained to capture human preferences across diverse perceptual dimensions. This enables models to learn complex and subjective visual criteria [xu2023imagereward, wu2023humanpreferencescorev2] in a data-free way. Despite its promise, applying RL to consistent image generation remains an unsolved challenge, primarily due to the absence of both (1) Suitable Reward Model and (2) Efficient RL Algorithm. On the reward side, existing reward models predominantly evaluate aesthetics and prompt alignment [ma2025hpsv3, wang2025unified, wu2025rewarddancerewardscalingvisual, wu2025editrewardhumanalignedrewardmodel], but the lack of explicit consistency evaluation makes them insufficient for consistent generation. On the optimization side, consistent generation is significantly more computationally demanding than single-image synthesis and requires carefully balancing consistency preservation and prompt fidelity. However, current RL algorithms struggle to achieve this trade-off efficiently.
@@
(2) PaCo-GRPO. Recent RL methods for image generation [liu2025flowgrpotrainingflowmatching, xue2025dancegrpounleashinggrpovisual, li2025mixgrpounlockingflowbasedgrpo] achieve strong results but remain computationally heavy, particularly for consistency tasks requiring high resolutions or multiple images per prompt. To improve efficiency and stability, we introduce PaCo-GRPO, which integrates two core strategies: (1) a resolution-decoupled training scheme that trains on low-resolution images, while preserving full-resolution generation at inference, effectively reducing computational cost without compromising final performance; and (2) a log-tamed multi-reward aggregation mechanism that balances rewards to prevent domination and ensure stable training. Together, these techniques significantly boost training efficiency and stability without compromising performance, and can be combined with prior optimizations [liu2025flowgrpotrainingflowmatching, li2025mixgrpounlockingflowbasedgrpo] to make consistent image generation both effective and computationally practical.
@@ 2.1 Consistent Image Generation
Consistent Image Generation aims to produce images coherent with given contexts, constraints, or attributes, encompassing tasks such as Text-to-ImageSet [tewel2024trainingfreeconsistenttexttoimagegeneration, hertz2024stylealignedimagegeneration, rahman2023make, jia2025settleonetexttoimagesetgeneration, song2025makeanythingharnessingdiffusiontransformers] and Image Editing [wang2025seededit30fasthighquality, wu2025editrewardhumanalignedrewardmodel, wu2025editrewardhumanalignedrewardmodel, luo2025editscoreunlockingonlinerl] While differing in formulation, both share the goal of generating images faithful to shared identities, styles, or visual conditions. Text-to-ImageSet methods rely on grouped or pairwise data with strong internal coherence [li2023photomaker, mao2024storyadaptertrainingfreeiterativeframework, liu2025onepromptonestory, huang2024consistentidportraitgenerationmultimodal, zhou2024storydiffusion, huang2024incontextloradiffusiontransformers], whereas image editing approaches [batifol2025flux, wu2025qwenimagetechnicalreport, chen2025adiee] learn controlled modifications from paired original-edited images, preserving fidelity while adjusting specific attributes.
@@ 2.2.1 Reward Models for Image Generation
Early works built reward models using existing image quality metrics or fine-tuned CLIP-based models on human preference datasets [wu2023human, wu2023human2, kirstain2023pickapicopendatasetuser, zhang2024learning, chen2025adiee, lin2024evaluating]. Recent studies adopt multimodal large language models (MLLMs) as backbones for reward modeling [ma2025hpsv3, wang2025unified, wu2025rewarddancerewardscalingvisual, wu2025editrewardhumanalignedrewardmodel, luo2025editscoreunlockingonlinerl, seedream2025seedream40nextgenerationmultimodal], demonstrating strong ability to understand complex human preferences and much greater flexibility across diverse evaluation tasks. However, these works most remain focused on general image-text alignment or aesthetic evaluation, largely ignoring the aspects of visual consistency.
In contrast, consistency-related preferences require pairwise image comparisons rather than simple image-text alignment or single-image quality assessment. Although models such as CLIP [radford2021learningtransferablevisualmodels] and DreamSim [fu2023dreamsim] can assess image similarity, they are not designed to capture the multifaceted nature of human perception of consistency. To bridge this gap, we propose a pairwise reward modeling framework that effectively captures human preferences for consistency across multiple aspects.
@@ 2.2.2 RL Algorithms for Image Generation
RL-style algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms, black2024training, fan2024reinforcement, gupta2025simple, miao2024training, zhao2025score], Direct Preference Optimization (DPO) [rafailov2024direct, wallace2024diffusion, yang2024using, liang2024step, yuan2024self, liu2024videodpo, zhang2024onlinevpo, furuta2024improving, liang2025aesthetic], and Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical, liu2025flowgrpotrainingflowmatching, xue2025dancegrpounleashinggrpovisual, li2025branchgrpostableefficientgrpo, he2025tempflowgrpotimingmattersgrpo, li2025mixgrpounlockingflowbasedgrpo, wang2025coefficientspreservingsamplingreinforcementlearning] have emerged as powerful paradigms to enhance generative models [peng2019advantage, fan2025online, wang2025prefgrpopairwisepreferencerewardbased, zheng2025diffusionnftonlinediffusionreinforcement, yeh2024training, tang2024tuning, song2023loss], and have demonstrated significant success in aligning text-to-image models with human preferences [prabhudesai2023aligning, clark2023directly, xu2023imagereward, prabhudesai2024video, lee2023aligning, dong2023raft].
Recent methods like Flow-GRPO [liu2025flowgrpotrainingflowmatching] and Dance-GRPO [xue2025dancegrpounleashinggrpovisual] introduce stochasticity to flow-matching models by converting ordinary differential equations (ODEs) into stochastic differential equations (SDEs), improving sampling diversity. However, RL-based image generation remains computationally expensive, with sampling being a key bottleneck. Techniques such as MixGRPO [liu2025flowgrpotrainingflowmatching, li2025mixgrpounlockingflowbasedgrpo] reduce training costs by mixing SDE and ODE sampling, but sampling overhead persists, particularly for consistent image generation tasks that handle multiple images.
To mitigate this, we propose PaCo-GRPO, a resolution-decoupled training strategy that samples low-resolution images during RL, significantly reducing computational demands. This approach complements existing methods like MixGRPO [li2025mixgrpounlockingflowbasedgrpo] and FlowGRPO-Fast [liu2025flowgrpotrainingflowmatching], offering additional efficiency gains when combined.
@@ 3.1 Paco-Dataset: Pairwise Consistency Dataset
Data Synthesis. Consistent generation involves multiple images sharing common elements while varying in other aspects, making data collection particularly challenging. To efficiently construct a large-scale dataset with diverse consistency patterns, we first generate 2,000 text prompts for Text-to-ImageSet synthesis using Deepseek-V3.1 [deepseekai2025deepseekv3technicalreport], then select 708 diverse prompts through graph-based diversification on text embeddings [radford2021learningtransferablevisualmodels]. Inspired by [huang2024incontextloradiffusiontransformers], we use FLUX.1-dev [blackforestlabs_flux1dev_2024] to generate $m\times n$ image grids with strong internal consistency. To enhance diversity and reduce cost, we generate four grids with different seeds per prompt and apply a sub-figure combinatorial pairing strategy, where we divide each grid into $m\times n$ subfigures and exhaustively pair them across grids from the same prompt. We empirically choose a $2\times 2$ grid setting as it offers the best trade-off between quality and efficiency. This process yields 33,984 unique ranking instances from 708 prompts and 2,832 images, substantially expanding dataset scale and diversity at minimal cost. Each instance comprises one reference image and four comparison candidates to be ranked by consistency, as illustrated in Fig.˜2 (top-left).
@@
Directly training on ranking data is computationally demanding and often requires non-trivial architectural changes [wu2025editrewardhumanalignedrewardmodel]. To retain both simplicity and generality, we convert rankings into pairwise comparisons (see Fig.˜2, top-right), preserving rich supervision while providing clear positive and negative pairs for consistency evaluation. To further improve diversity and balance, we incorporate 5,695 manually verified consistent pairs from ShareGPT-4o-Image [chen2025sharegpt4oimagealigningmultimodalmodels]. In total, the dataset contains 54,624 annotated image pairs (27,599 consistent and 27,025 inconsistent), each labeled through human evaluation and augmented with CoT reasoning annotations generated by GPT-5 [OpenAI2025IntroducingGPT5]. These reasoning annotations enhance dataset interpretability and help mitigate overfitting during Paco-Reward training.
@@ 3.2 PaCo-Reward: Pairwise Reward Modeling
Motivation. Consistency evaluation inherently requires comparing multiple images to assess their alignment on shared attributes. Recent works [luo2025editscoreunlockingonlinerl, wu2025editrewardhumanalignedrewardmodel] demonstrate that VLM-based reward models are effective for such multi-image, text-conditioned tasks. However, existing approaches have key limitations. Some introduce extra regression heads to output scalar rewards, which mismatch with the next-token prediction nature of VLMs. Others rely on long CoT reasoning to infer reward scores, causing high computational overhead during RL training.
Reward Modeling. Inspired by prior works [jia2025settleonetexttoimagesetgeneration, xue2025dancegrpounleashinggrpovisual], we propose PaCo-Reward, a pairwise reward modeling framework that simplifies consistency learning. Rather than modeling full rankings, it learns relative preferences from pairwise comparisons to provide rich and efficient supervision. To align with the autoregressive nature of VLMs, PaCo-Reward further reformulates consistency evaluation as a generative task, predicting the likelihood of “Yes” or “No” to indicate the consistency between two images. This formulation naturally fits the autoregressive next-token prediction paradigm of VLMs and can be further enhanced with CoT reasoning to improve interpretability and robustness, as illustrated in Fig.˜2 (top-right). During inference, the predicted probability of “Yes” serves as the consistency score between two images, thus facilitating efficient pairwise consistency evaluation for RL training. Moreover, human preference rankings can be inferred from the consistency scores of candidate images relative to the reference image, as shown in Fig.˜2 (bottom-left).
@@
$\mathcal{L}_{\text{PaCo}}=-\left[\alpha\log p(y_{0}\mid I)+(1-\alpha)\sum_{i=1}^{n-1}\log p(y_{i}\mid I)\right],$
@@ 4.1 Preliminaries
Recently, FlowGRPO [liu2025flowgrpotrainingflowmatching] and DanceGRPO [xue2025dancegrpounleashinggrpovisual] have successfully adapted Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] to image generation models. These methods introduce stochasticity into flow-matching models by converting deterministic ODEs into SDEs. The corresponding SDE update rule is:
@@
Consistent image generation tasks often involve multiple images and require the simultaneous optimization of several reward signals, such as visual consistency and prompt alignment. This presents significant challenges in two aspects: (i) RL efficiency: optimizing large or multiple images substantially increases computational cost. Prior work [li2025mixgrpounlockingflowbasedgrpo, he2025tempflowgrpotimingmattersgrpo, liu2025flowgrpotrainingflowmatching] demonstrates that training can be reduced to 1-2 steps without compromising performance, yet sampling efficiency, which remains the primary computational bottleneck, is largely unaddressed. (ii) RL stability: multi-reward optimization is prone to reward domination [roijers2013survey, hayes2021practical], where one signal can dominate others, resulting in suboptimal outcomes or destabilized training.
@@ 4.2 PaCo-GRPO Strategies
Resolution-decoupled Training. In Text-to-ImageSet generation, models often produce high-resolution outputs containing multiple sub-figures, each matching the standard resolution of text-to-image tasks. This requirement dramatically increases computational cost, scaling quadratically with image resolution in Transformer-based architectures [vaswani2017attention]. Such scaling makes RL particularly expensive.
FlowGRPO [liu2025flowgrpotrainingflowmatching] has observed that even low-quality images generated with fewer denoising steps can still provide effective reward signals for RL optimization. Building on this insight, we introduce a resolution-decoupled training strategy to improve efficiency. While the model generates high-resolution ($h\times w$) images during inference and evaluation, it produces lower-resolution images (e.g., $\frac{h}{2}\times\frac{w}{2}$) during training for reward computation and optimization. This approach substantially reduces the overhead of both sampling and training, thereby accelerating RL optimization (see Fig.˜6). It also integrates seamlessly with existing strategies [li2025mixgrpounlockingflowbasedgrpo, liu2025flowgrpotrainingflowmatching, he2025tempflowgrpotimingmattersgrpo] for further efficiency gains.
Log-tamed Multi-reward Aggregation During RL training [schulman2017trustregionpolicyoptimization, schulman2017proximalpolicyoptimizationalgorithms, rafailov2024direct, shao2024deepseekmathpushinglimitsmathematical], multiple reward signals are often combined to guide models toward desired behaviors. For consistent image generation, both visual consistency and prompt alignment are essential for ensuring high-quality outputs. Given $N$ input conditions $\{\bm{c}_{i}\}_{i=1}^{N}$, the model generates a group of $G$ samples $\{\bm{x}_{i}^{j}\}_{j=1}^{G}$ for each condition $\bm{c}_{i}$. Each reward model $R^{k}$ outputs a scalar score $R^{k}(\bm{x}_{i}^{j},\bm{c}_{i})$, and the aggregated reward and advantage are computed as
@@
where $w_{k}$ is the weight for the $k$-th reward. However, this naive aggregation often suffers from reward domination [roijers2013survey, hayes2021practical], where one reward dominates optimization, leading to suboptimal or even degenerate results.
This issue can be mitigated by manual weight tuning, but the process is labor-intensive and lacks generality. To address this, we propose a log-tamed aggregation strategy, which adaptively compresses rewards with large fluctuations, thus preventing them from overwhelming the overall optimization. Specifically, we first compute the coefficient of variation [wikipedia_coefficient_of_variation] for the $k$-th reward, denoted as $h^{k}$:

### v2

@@ Abstract
Consistent image generation requires faithfully preserving identities, styles, and logical coherence across multiple images, which is essential for applications such as storytelling and character design. Supervised training approaches struggle with this task due to the lack of large-scale datasets capturing visual consistency and the complexity of modeling human perceptual preferences. In this paper, we argue that reinforcement learning (RL) offers a promising alternative by enabling models to learn complex and subjective visual criteria in a data-free manner. To achieve this, we introduce PaCo-RL, a comprehensive framework that combines a specialized consistency reward model with an efficient RL algorithm. The first component, PaCo-Reward, is a pairwise consistency evaluator trained on a large-scale dataset constructed via automated sub-figure pairing. It evaluates consistency through a generative, autoregressive scoring mechanism enhanced by task-aware instructions and CoT reasons. The second component, PaCo-GRPO, leverages a novel resolution-decoupled optimization strategy to substantially reduce RL cost, alongside a log-tamed multi-reward aggregation mechanism that ensures balanced and stable reward optimization. Extensive experiments across the two representative subtasks show that PaCo-Reward significantly improves alignment with human perceptions of visual consistency, and PaCo-GRPO achieves state-of-the-art consistency performance with improved training efficiency and stability. Together, these results highlight the promise of PaCo-RL as a practical and scalable solution for consistent image generation. Project page.
@@ 1 Introduction
Image generation has made remarkable progress in producing high-quality and diverse images from textual prompts [56, 3, 12, 46, 78, 25, 85, 23, 69, 70]. However, achieving consistent image generation remains both challenging and crucial for various applications such as storytelling, advertising, and character creation [42, 65, 5, 97, 39, 22]. In this work, we study consistent image generation through two representative tasks: Image Editing [40] and Text-to-ImageSet [24], where the objective is to either align generated images with given references (Fig.˜1, left) or produce a coherent set of images from a single prompt (Fig.˜1, right). This is a challenging problem for supervised training approaches, due to the lack of large-scale datasets for visual consistency and the complexity of modeling human perceptions of consistency [83, 26, 93, 7, 36].
Recently, reinforcement learning (RL) has demonstrated strong potential in advancing image generation models [37, 88, 95], offering a promising pathway to address the above challenges. Instead of relying on explicit supervision from carefully curated datasets, RL algorithms optimize generative models using feedback from reward models (RMs), which are often trained to capture human preferences across diverse perceptual dimensions. This enables models to learn complex and subjective visual criteria [87, 83] in a data-free way. Despite its promise, applying RL to consistent image generation remains an unsolved challenge, primarily due to the absence of both (1) Suitable Reward Model and (2) Efficient RL Algorithm. On the reward side, existing reward models predominantly evaluate aesthetics and prompt alignment [41, 76, 80, 81], but the lack of explicit consistency evaluation makes them insufficient for consistent generation. On the optimization side, consistent generation is significantly more computationally demanding than single-image synthesis and requires carefully balancing consistency preservation and prompt fidelity. However, current RL algorithms struggle to achieve this trade-off efficiently.
@@
(2) PaCo-GRPO. Recent RL methods for image generation [37, 88, 31] achieve strong results but remain computationally heavy, particularly for consistency tasks requiring high resolutions or multiple images per prompt. To improve efficiency and stability, we introduce PaCo-GRPO, which integrates two core strategies: (1) a resolution-decoupled training scheme that trains on low-resolution images, while preserving full-resolution generation at inference, effectively reducing computational cost without compromising final performance; and (2) a log-tamed multi-reward aggregation mechanism that balances rewards to prevent domination and ensure stable training. Together, these techniques significantly boost training efficiency and stability without compromising performance, and can be combined with prior optimizations [37, 31] to make consistent image generation both effective and computationally practical.
@@ 2.1 Consistent Image Generation
Consistent Image Generation aims to produce images coherent with given contexts, constraints, or attributes, encompassing tasks such as Text-to-ImageSet [65, 20, 54, 24, 62] and Image Editing [73, 81, 81, 40] While differing in formulation, both share the goal of generating images faithful to shared identities, styles, or visual conditions. Text-to-ImageSet methods rely on grouped or pairwise data with strong internal coherence [33, 42, 39, 21, 97, 22], whereas image editing approaches [2, 78, 7, 71, 72] learn controlled modifications from paired original-edited images, preserving fidelity while adjusting specific attributes.
@@ 2.2.1 Reward Models for Image Generation
Early works built reward models using existing image quality metrics or fine-tuned CLIP-based models on human preference datasets [84, 82, 26, 93, 7, 36]. Recent studies adopt multimodal large language models (MLLMs) as backbones for reward modeling [41, 76, 80, 81, 40, 59], demonstrating strong ability to understand complex human preferences and much greater flexibility across diverse evaluation tasks. However, these works most remain focused on general image-text alignment or aesthetic evaluation, largely ignoring the aspects of visual consistency.
In contrast, consistency-related preferences require pairwise image comparisons rather than simple image-text alignment or single-image quality assessment. Although models such as CLIP [52] and DreamSim [15] can assess image similarity, they are not designed to capture the multifaceted nature of human perception of consistency. To bridge this gap, we propose a pairwise reward modeling framework that effectively captures human preferences for consistency across multiple aspects.
@@ 2.2.2 RL Algorithms for Image Generation
RL-style algorithms such as Proximal Policy Optimization (PPO) [58, 4, 14, 17, 43, 94], Direct Preference Optimization (DPO) [53, 67, 89, 35, 91, 38, 92, 16, 34], and Group Relative Policy Optimization (GRPO) [60, 37, 88, 32, 19, 31, 68] have emerged as powerful paradigms to enhance generative models [48, 13, 75, 95, 90, 63, 61], and have demonstrated significant success in aligning text-to-image models with human preferences [49, 8, 87, 50, 30, 11].
Recent methods like Flow-GRPO [37] and Dance-GRPO [88] introduce stochasticity to flow-matching models by converting ordinary differential equations (ODEs) into stochastic differential equations (SDEs), improving sampling diversity. However, RL-based image generation remains computationally expensive, with sampling being a key bottleneck. Techniques such as MixGRPO [37, 31] reduce training costs by mixing SDE and ODE sampling, but sampling overhead persists, particularly for consistent image generation tasks that handle multiple images.
To mitigate this, we propose PaCo-GRPO, a resolution-decoupled training strategy that samples low-resolution images during RL, significantly reducing computational demands. This approach complements existing methods like MixGRPO [31] and FlowGRPO-Fast [37], offering additional efficiency gains when combined.
@@ 3.1 Paco-Dataset: Pairwise Consistency Dataset
Data Synthesis. Consistent generation involves multiple images sharing common elements while varying in other aspects, making data collection particularly challenging. To efficiently construct a large-scale dataset with diverse consistency patterns, we first generate 2,000 text prompts for Text-to-ImageSet synthesis using Deepseek-V3.1 [10], then select 708 diverse prompts through graph-based diversification on text embeddings [52]. Inspired by [22], we use FLUX.1-dev [29] to generate $m\times n$ image grids with strong internal consistency. To enhance diversity and reduce cost, we generate four grids with different seeds per prompt and apply a sub-figure combinatorial pairing strategy, where we divide each grid into $m\times n$ subfigures and exhaustively pair them across grids from the same prompt. We empirically choose a $2\times 2$ grid setting as it offers the best trade-off between quality and efficiency. This process yields 33,984 unique ranking instances from 708 prompts and 2,832 images, substantially expanding dataset scale and diversity at minimal cost. Each instance comprises one reference image and four comparison candidates to be ranked by consistency, as illustrated in Fig.˜2 (top-left).
@@
Directly training on ranking data is computationally demanding and often requires non-trivial architectural changes [81]. To retain both simplicity and generality, we convert rankings into pairwise comparisons (see Fig.˜2, top-right), preserving rich supervision while providing clear positive and negative pairs for consistency evaluation. To further improve diversity and balance, we incorporate 5,695 manually verified consistent pairs from ShareGPT-4o-Image [6]. In total, the dataset contains 54,624 annotated image pairs (27,599 consistent and 27,025 inconsistent), each labeled through human evaluation and augmented with CoT reasoning annotations generated by GPT-5 [47]. These reasoning annotations enhance dataset interpretability and help mitigate overfitting during Paco-Reward training.
@@ 3.2 PaCo-Reward: Pairwise Reward Modeling
Motivation. Consistency evaluation inherently requires comparing multiple images to assess their alignment on shared attributes. Recent works [40, 81] demonstrate that VLM-based reward models are effective for such multi-image, text-conditioned tasks. However, existing approaches have key limitations. Some introduce extra regression heads to output scalar rewards, which mismatch with the next-token prediction nature of VLMs. Others rely on long CoT reasoning to infer reward scores, causing high computational overhead during RL training.
Reward Modeling. Inspired by prior works [24, 88], we propose PaCo-Reward, a pairwise reward modeling framework that simplifies consistency learning. Rather than modeling full rankings, it learns relative preferences from pairwise comparisons to provide rich and efficient supervision. To align with the autoregressive nature of VLMs, PaCo-Reward further reformulates consistency evaluation as a generative task, predicting the likelihood of “Yes” or “No” to indicate the consistency between two images. This formulation naturally fits the autoregressive next-token prediction paradigm of VLMs and can be further enhanced with CoT reasoning to improve interpretability and robustness, as illustrated in Fig.˜2 (top-right). During inference, the predicted probability of “Yes” serves as the consistency score between two images, thus facilitating efficient pairwise consistency evaluation for RL training. Moreover, human preference rankings can be inferred from the consistency scores of candidate images relative to the reference image, as shown in Fig.˜2 (bottom-left).
@@
$\mathcal{L}_{\text{PaCo}}=-\left[\alpha\log p(y_{0}\mid I)+\frac{(1-\alpha)}{n-1}\sum_{i=1}^{n-1}\log p(y_{i}\mid I)\right],$
@@ 4.1 Preliminaries
Recently, FlowGRPO [37] and DanceGRPO [88] have successfully adapted Group Relative Policy Optimization (GRPO) [60] to image generation models. These methods introduce stochasticity into flow-matching models by converting deterministic ODEs into SDEs. The corresponding SDE update rule is:
@@
Consistent image generation tasks often involve multiple images and require the simultaneous optimization of several reward signals, such as visual consistency and prompt alignment. This presents significant challenges in two aspects: (i) RL efficiency: optimizing large or multiple images substantially increases computational cost. Prior work [31, 19, 37] demonstrates that training can be reduced to 1-2 steps without compromising performance, yet sampling efficiency, which remains the primary computational bottleneck, is largely unaddressed. (ii) RL stability: multi-reward optimization is prone to reward domination [55, 18], where one signal can dominate others, resulting in suboptimal outcomes or destabilized training.
@@ 4.2 PaCo-GRPO Strategies
Resolution-decoupled Training. In Text-to-ImageSet generation, models often produce high-resolution outputs containing multiple sub-figures, each matching the standard resolution of text-to-image tasks. This requirement dramatically increases computational cost, scaling quadratically with image resolution in Transformer-based architectures [66]. Such scaling makes RL particularly expensive.
FlowGRPO [37] has observed that even low-quality images generated with fewer denoising steps can still provide effective reward signals for RL optimization. Building on this insight, we introduce a resolution-decoupled training strategy to improve efficiency. While the model generates high-resolution ($h\times w$) images during inference and evaluation, it produces lower-resolution images (e.g., $\frac{h}{2}\times\frac{w}{2}$) during training for reward computation and optimization. This approach substantially reduces the overhead of both sampling and training, thereby accelerating RL optimization (see Fig.˜6). It also integrates seamlessly with existing strategies [31, 37, 19] for further efficiency gains.
Log-tamed Multi-reward Aggregation During RL training [57, 58, 53, 60], multiple reward signals are often combined to guide models toward desired behaviors. For consistent image generation, both visual consistency and prompt alignment are essential for ensuring high-quality outputs. Given $N$ input conditions $\{\bm{c}_{i}\}_{i=1}^{N}$, the model generates a group of $G$ samples $\{\bm{x}_{i}^{j}\}_{j=1}^{G}$ for each condition $\bm{c}_{i}$. Each reward model $R^{k}$ outputs a scalar score $R^{k}(\bm{x}_{i}^{j},\bm{c}_{i})$, and the aggregated reward and advantage are computed as
@@
where $w_{k}$ is the weight for the $k$-th reward. However, this naive aggregation often suffers from reward domination [55, 18], where one reward dominates optimization, leading to suboptimal or even degenerate results.
This issue can be mitigated by manual weight tuning, but the process is labor-intensive and lacks generality. To address this, we propose a log-tamed aggregation strategy, which adaptively compresses rewards with large fluctuations, thus preventing them from overwhelming the overall optimization. Specifically, we first compute the coefficient of variation [77] for the $k$-th reward, denoted as $h^{k}$: